# A First Look at Package-to-Group Mechanism: An Empirical Study of the Linux Distributions

### v1

@@ Abstract
Reusing third-party software packages is a common practice in software development. As the scale and complexity of open-source software (OSS) projects continue to grow (e.g., Linux distributions), the number of reused third-party packages has significantly increased. Therefore, maintaining effective package management is critical for developing and evolving OSS projects. To achieve this, a package-to-group mechanism (P2G) is employed to enable unified installation, uninstallation, and updates of multiple packages at once. To better understand this mechanism, this paper takes Linux distributions as a case study and presents an empirical study focusing on its application trends, evolutionary patterns, group quality, and developer tendencies. By analyzing 11,746 groups and 193,548 packages from 89 versions of 5 popular Linux distributions and conducting questionnaire surveys with Linux practitioners and researchers, we derive several key insights. Our findings show that P2G is increasingly being adopted, particularly in popular Linux distributions. P2G follows six evolutionary patterns (e.g., splitting and merging groups). Interestingly, packages no longer managed through P2G are more likely to remain in Linux distributions rather than being directly removed. To assess the effectiveness of P2G, we propose a metric called GValue to evaluate the quality of groups and identify issues such as inadequate group descriptions and insufficient group sizes. We also summarize five types of packages that tend to adopt P2G, including graphical desktops, networks, etc. To the best of our knowledge, this is the first study focusing on the P2G mechanisms. We expect our study can assist in the efficient management of packages and reduce the burden on practitioners in rapidly growing Linux distributions and other open-source software projects.
@@ I Introduction
The development of modern software systems heavily relies on code reuse, often achieved through the use of third-party packages [1]. Utilizing packages to build software can reduce development costs and enhance productivity [2][3]. Due to these advantages, third-party packages are also widely used in open-source software (OSS) projects [4]. However, with the rapid rise in popularity of OSS projects due to their low cost and high quality [5][6], their scale and complexity have also significantly increased [7][8], resulting in a sharp rise in the reuse of third-party packages. For instance, CentOS version 7 contains 14,479 packages, while version 3.7 released in 2006 has only 1,275 packages [9]. This growth presents a critical challenge: mazhenaging packages effectively to ensure the ongoing development, continuous evolution, and long-term success of OSS projects.
@@
In our literature review, we found that there has been no prior research on the P2G mechanism. Previous studies related to package management mainly focus on dependency analysis [15][16], freshness evaluation [17][18], and malicious detection [19][20]. Additionally, some studies have discussed package layering [21] and classification [22]. Their core concepts are similar to P2G, as they all involve dividing packages into different sets (i.e., groups, layers, or categories). Compared to P2G, these approaches use broader, more coarse-grained divisions. This leads to large groups with too many packages, making it difficult to manage or update packages when a particular feature of the OSS project changes.
@@
In this paper, considering Linux is a classic example of OSS projects due to its widespread adoption and vast ecosystem of packages [23], we take Linux distributions as a case study to investigate the P2G mechanism in an evolutionary context. We start by understanding the application trends of this mechanism and analyzing the evolutionary patterns across previous versions. Then we propose a metric called GValue to evaluate the quality and identify issues in current versions. Finally, we explore the developer tendencies to support P2G in future versions. In particular, this paper addresses the following research questions (RQs):
@@
RQ1: How prevalent is the use of the P2G mechanism? We carefully selected 5 Linux distributions and identified groups, packages within those groups, and the total number of packages in each distribution. We analyzed the trend of adopting this mechanism over versions, the distributions of packages within the P2G, and the correlation between P2G adoption and distribution popularity. Our analysis shows that while P2G is becoming increasingly common, especially in popular distributions, it still accounts for only a small proportion of total packages across Linux distributions.
@@
RQ2: What are the evolutionary patterns in the P2G mechanism? We address this from two aspects: a content-based summary of the change patterns and a quantitative analysis of package flows. We manually analyzed the addition and removal of groups across 89 consecutive versions of the 5 selected Linux distributions to summarize the change patterns of P2G and compute the number of packages affected in each scenario. Then, we developed a taxonomy of these change patterns and found that packages no longer adopting the P2G are more likely to remain in Linux distributions.
@@
RQ3: What is the quality of the current groups in the P2G mechanism? To evaluate group quality, we established a multi-dimensional feature method called GValue. To validate its effectiveness, we randomly selected 5 groups to design a questionnaire survey for manual scoring, followed by a correlation analysis with GValue. Our results show that 16% of the groups exhibit poor quality (i.e., GValue score below 0.2) across Linux distributions. Additionally, we identified several existing issues with P2G adoption from an analysis of 413 groups in the latest version distributions.
@@
RQ4: What types of packages do developers tend to adopt the P2G mechanism? We performed a topic analysis on group description and identified 5 categories (e.g., Graphical Desktop, etc) that are more likely to adopt this mechanism. Furthermore, we extracted key information from package descriptions and compared them between packages with the P2G and without the P2G. This analysis provides valuable insights for developers and maintainers in determining whether a package should adopt this mechanism.
@@
  • We provide a comprehensive analysis of the P2G mechanism covering its application trends, evolutionary patterns, group quality, and developer tendencies.
  • We summarize various evolutionary patterns in groups, offering insights that can guide OSS project maintainers in adopting the P2G in future updates.
@@
  • We discover 5 categories of packages where developers frequently adopt the P2G, offering valuable guidance for Linux maintainers in deciding whether to implement P2G for specific packages.
@@ II-A Related Work
Various theories and strategies have been proposed to reduce the workload of package management and enhance the productivity of OSS projects [27]. Wang et al. [28] introduced a graphical method to establish dependency relations within a Linux distribution and analyzed the complex graph of related attributes. Sousa et al. [29] divided packages into communities based on their dependencies, recommending that packages in the same community be maintained by the same team according to the strength of their relationships. Legay et al. [18] evaluated and compared the freshness of 890 packages across six mainstream Linux distributions. Halder et al. [19] proposed a metadata-based model for malicious package detection. Thus, previous studies mainly focus on package dependency analysis, freshness evaluation, and malicious detection. There is a lack of in-depth research on the package-to-group mechanism in OSS.
Though no prior studies directly focus on the P2G mechanism, some related works exist  [21] [22], all of which classify multiple packages into distinct sets. Tan et al.  [21] suggested that a clear package hierarchy can facilitate the rapid construction and maintenance of Linux distributions. They proposed an efficient and accurate package layering algorithm that divides packages into four layers. However, their approach categorizes packages into only four levels, resulting in an excessive number of packages within the same level. While this approach aids in building integrated Linux and retrieving specific packages, its overall effectiveness is limited. Jing et al. [22] argue that package categorization can guide Linux developers and manufacturers in selecting components and provide guidance during the construction process. They created a categorized dataset of packages and outlined selection boundaries for component architecture, offering useful insights for the construction process. However, this work only focuses on a single Linux distribution, and the number of categories remains too few to handle the complexity of OSS projects.
To further advance these works, this paper focuses on the P2G mechanism in Linux distributions, investigating its application trends, evolutionary patterns, group quality, and developer tendencies. This work aims to assist developers and maintainers in the open-source community to efficiently manage packages by adopting the P2G mechanism.
@@ II-B Package-to-group Mechanism
The package-to-group mechanism bundles multiple packages with similar purposes into a group, as shown in Figure 1. This mechanism primarily simplifies the installation and management, enhancing efficiency for developers and maintainers. Different Linux distributions implement this mechanism in different ways. For instance, Debian and Debian-based systems (e.g., Ubuntu), use metapackages [30], which define groups of related packages based on their purposes. This allows maintainers to install a complete set of packages for a specific task, such as setting up a desktop or development environment, in one step. In Red Hat and its derivatives (e.g., Fedora and CentOS), this mechanism is based on the concept of groups. These groups can be defined by a configuration file, which lists relevant groups and their associated packages. Maintainers can easily install or uninstall groups, making it easier to manage large sets of related packages.
The P2G mechanism not only helps maintainers more easily locate and install the packages they need but also facilitates system maintenance. Additionally, this mechanism can simplify security updates and upgrades, as OSS projects can be targeted at specific groups rather than updating each package individually. This is particularly important in managing servers and large enterprise environments.
@@ III Overview
In this section, we present the design of our study and provide an overview of the Linux distributions we analyzed.
@@ III-A Study Design
Figure 2 provides an overview of our study, which focuses on five Linux distributions (Section III-B). Our study follows the evolutionary process and can be divided into three phases (i.e., the past, the present, and the future). In the past phase, we conduct quantitative analysis and change analysis on historical versions to address questions regarding P2G’s application (RQ1) and evolution (RQ2). In the present phase, we assess the latest version’s quality (RQ3) by proposing metrics, evaluating groups, and summarizing issues. In the future phase, we conduct a topic analysis for group descriptions and key information extraction from package descriptions to explore developer tendencies (RQ4). The practical implications of our research include providing current Linux maintainers with update recommendations based on evolutionary patterns and quality issues and helping Linux developers determine whether to adopt the P2G based on developer tendencies.
@@
III-B Studied Linux Distributions
There are many package management tools for Linux, such as RPM for Red Had-based distributions and DEB for Debian-based distributions. Considering RPM has become one of the recognized standards in Linux distributions [12], we rank Linux distributions under RPM based on developers’ interests [31] and select the top three from them, including Fedora, CentOS, and OpenEuler. These Linux are widely used general-purpose distributions. To provide broader coverage of Linux across different scenarios, we also include two distributions tailored for specific cases: Anolis [32] and OpenCloudOS [33]. Anolis focuses on enterprise-level applications and is a popular choice for data centers. OpenCloudOS is designed for cloud computing environments and is suitable for cloud service providers. Table I shows the studied Linux distributions and their versions. We obtained package and group information for each distribution from official or widely available mirrors.
@@ IV-A Methodology
We answer this RQ from three aspects: 1) the trend in the number of groups, the number of packages adopting P2G, and the total number of packages; 2) the distribution of packages adopting P2G within a Linux distribution; 3) the relationship between the adoption of P2G and distribution popularity. The analysis is based on the selected Linux distributions and release versions in Section III-B. To decide whether a package adopts the P2G in a specific release version, we consider whether it is included in a group in that version.
@@ IV-B Results
Application Trends. Figure 3 shows the number of groups, packages adopting P2G, and total packages across historical release versions. ❶ We observe that the number of groups remains relatively stable whether in a complex Linux distribution (e.g., Fedora) or a relatively simplified distribution (e.g., OpenEuler). This suggests that Linux distributions have a relatively fixed group structure. ❷ However, most Linux distributions show an increase in the number of packages adopting P2G, especially in more innovative and developer-oriented distributions (e.g., Fedora). In contrast, in highly stable scenarios (e.g., enterprise or server environments), Linux distributions (e.g., openEuler) show less variation, as they prioritize reliability and compatibility over frequent changes. ❸ The number of total packages continues to increase with version updates. This reflects how Linux distributions continuously add new packages to provide more features and support in response to enhanced functionality and growing user requirements. For example in Fedora, the number of packages increased nearly tenfold, from 11,410 to 103,176.
@@
Distribution of packages adopting P2G. Figure 4 shows violin plots with the distribution of packages using P2G and the radio of packages adopting P2G. ❶ The number of packages adopting P2G varies significantly across distributions. Fedora has a median of 3,544 P2G-adopting packages and openEuler has only 776, which is nearly one-fifth of Fedora’s. ❷ The distribution of package numbers across versions also varies among Linux distributions. Anolis shows significant fluctuations, ranging from 0 to 4,000, whereas openEuler remains within a more stable range of 700 to 800, indicating minimal volatility. ❸ The number and radio of packages using P2G are not consistent. Fedora has the highest number of grouped packages (median of 3,544), but a lower ratio of 0.27. In contrast, CentOS has fewer grouped packages (a median of 1,297), yet it has the highest ratio at 0.47.
@@
Distribution of radio per version
@@
Relationship between the popularity of Linux distributions and the radio of packages adopting P2G. As shown in Figure 5, the radio shows a downward trend, except for Anolis. It suggests that a Linux distribution’s popularity is correlated with the application of P2G. To validate this assumption, we perform a Spearman correlation analysis between the radio of packages adopting P2G and the number of stars a Linux distribution has. However, the results (p-value=0.6, rho=0.4) indicate a weak positive correlation and the high p-value suggests that this correlation is statistically insignificant. Thus, it shows that the factors related to a Linux distribution’s popularity are complex. The anomaly observed in Anolis may be due to its inheritance of the CentOS ecosystem, resulting in a higher ratio despite its relatively low popularity.
Figure 5: Radio of the packages applied P2G mechanism. All the Linux distributions are sorted in descending order of popularity (#stars).

### v2

@@ Abstract
Reusing third-party software packages is a common practice in software development. As the scale and complexity of open-source software (OSS) projects continue to grow (e.g., Linux distributions), the number of reused third-party packages has significantly increased. Therefore, maintaining effective package management is critical for developing and evolving OSS projects. To achieve this, a package-to-group mechanism (P2G) is employed to enable unified installation, uninstallation, and updates of multiple packages at once. To better understand this mechanism, this paper takes Linux distributions as a case study and presents an empirical study focusing on its application trends, evolution patterns, group quality, and group tendency. By analyzing 11,746 groups and 193,548 packages from 89 versions of 5 popular Linux distributions and conducting questionnaire surveys with Linux practitioners and researchers, we derive several key insights. Our findings show that P2G is increasingly being adopted, particularly in popular Linux distributions. P2G follows six evolutionary patterns (e.g., splitting and merging groups). Interestingly, packages no longer managed through P2G are more likely to remain in Linux distributions rather than being directly removed. Besides, we propose a metric called GValue to evaluate the quality of groups and identify issues such as inadequate group descriptions and insufficient group sizes. We also summarize five types of packages that tend to adopt P2G, including graphical desktops and networks. To the best of our knowledge, this is the first study focusing on the P2G mechanisms. We expect our study can assist in the efficient management of packages and reduce the burden on practitioners in rapidly growing Linux distributions and other open-source software projects.
@@ I Introduction
The development of modern software systems heavily relies on code reuse, often achieved through the use of third-party packages [1]. Utilizing packages to build software can reduce development costs and enhance productivity [2][3]. Due to these advantages, third-party packages are also widely used in open-source software (OSS) projects [4]. However, with the rapid rise in popularity of OSS projects due to their low cost and high quality [5][6], their scale and complexity have also significantly increased [7][8], resulting in a sharp rise in the reuse of third-party packages. For instance, CentOS version 7 contains 14,479 packages, while version 3.7 released in 2006 has only 1,275 packages [9]. This growth presents a critical challenge: effectively managing packages to ensure the ongoing development, continuous evolution, and long-term success of OSS projects.
@@
In our literature review, we found that there has been no prior research on the P2G mechanism. Previous studies related to package management mainly focus on dependency analysis [15][16], freshness evaluation [17][18], and malicious detection [19][20]. Additionally, some studies have discussed package layering [21] and classification [22]. Their core concepts are similar to P2G, as they all involve dividing packages into different sets (i.e., groups, layers, or categories). Compared to P2G, these approaches employ broader, more coarse-grained divisions. This leads to large groups containing an excessive number of packages, which causes inconveniences in managing or updating packages when a particular feature of the OSS project changes.
@@
In this paper, we take Linux distributions as a case study to investigate the P2G mechanism in an evolution context. Linux distributions are chosen due to their representativeness for OSS projects and their widespread adoption and vast ecosystem of packages [23]. Specifically, we start by understanding the application trends of this mechanism and analyzing the evolutionary patterns across previous versions. Then we propose a metric called GValue to evaluate the quality and identify issues in current versions. Finally, we explore the group tendency to support P2G in future versions. In particular, this paper addresses the following research questions (RQs):
@@
RQ1: How common is the use of the P2G mechanism? We carefully select 5 Linux distributions and identified groups, packages within those groups, and the total number of packages in each distribution. We analyze the trend of adopting this mechanism over versions, the distributions of packages within P2G, and the correlation between P2G adoption and distribution popularity. Our analysis shows that while P2G is becoming increasingly common, especially in popular distributions, it still accounts for only a small proportion of total packages across Linux distributions.
@@
RQ2: What are the evolutionary patterns in the P2G mechanism? We address this from two perspectives: a content-based summary of change patterns and a quantitative analysis of package flows. Specifically, we manually analyze the addition and removal of groups across 89 consecutive versions of five Linux distributions to summarize the P2G change patterns and calculate the number of affected packages in each flow scenario. Finally, we build a taxonomy of these change patterns and find that packages no longer adopting P2G are more likely to remain in Linux distributions.
@@
RQ3: What is the quality of the current groups in the P2G mechanism? To evaluate group quality, we establish a multi-dimensional feature method called GValue. To validate its effectiveness, we randomly select 5 groups to design a questionnaire survey for manual scoring, followed by a correlation analysis [37] with GValue. Our results show that 16% of the groups exhibit poor quality (i.e., GValue score below 0.2) across Linux distributions. Additionally, we identify several existing issues on P2G adoption from an analysis of 413 groups in the latest version.
@@
RQ4: What types of packages do developers tend to adopt the P2G mechanism? We conduct a topic analysis on group description and identify 5 categories (e.g., Graphical Desktop) that are more likely to adopt P2G. Furthermore, we extract key information from package descriptions and compare them between packages within and without P2G. This analysis provides insights for developers and maintainers in deciding whether a package should adopt P2G.
@@
  • We provide a comprehensive analysis of the P2G mechanism covering its application trends, evolutionary patterns, group quality, and group tendency.
  • We summarize various evolutionary patterns in groups, offering insights that guide OSS project maintainers in adopting P2G in future updates.
@@
  • We discover 5 categories of packages where developers frequently adopt P2G, offering valuable guidance for Linux maintainers in deciding whether to adopt P2G for a specific package.
@@ II-A Related Work
Various theories and strategies have been proposed to reduce the workload of package management and enhance the productivity of OSS projects [27]. Wang et al. [28] introduced a graphical method to establish dependency relations within a Linux distribution and analyzed the complex graph of related attributes. Sousa et al. [29] divided packages into communities based on their dependencies, recommending that the same team maintain packages in the same community according to the strength of their relationships. Legay et al. [18] evaluated and compared the freshness of 890 packages across six mainstream Linux distributions. Halder et al. [19] proposed a metadata-based model for malicious package detection. Thus, previous studies mainly focus on package dependency analysis, freshness evaluation, and malicious detection. There is a lack of in-depth research on the package-to-group mechanism in OSS projects.
Although no prior studies directly focus on the P2G mechanism, some related works exist  [21] [22]. They all involve classifying multiple packages into distinct sets. Tan et al.  [21] suggested that a clear package hierarchy can facilitate the rapid construction and maintenance of Linux distributions. They proposed an efficient and accurate package layering algorithm that divides packages into four layers. However, their approach categorizes packages into only four levels, resulting in an excessive number of packages within the same level. While this approach aids in building integrated Linux and retrieving specific packages, its overall effectiveness is limited. Jing et al. [22] argue that package categorization can guide Linux developers in selecting components and provide guidance during the construction process. They created a categorized dataset of packages and outlined selection boundaries for component architecture, offering useful insights for the construction process. However, this work only focuses on a single Linux distribution, and the number of categories remains too few to handle the complexity of OSS projects.
To further advance these works, this paper focuses on the P2G mechanism in Linux distributions, investigating its application trends, evolutionary patterns, group quality, and group tendency. This work aims to assist developers and maintainers in the OSS projects efficiently manage packages by adopting the P2G mechanism.
@@ II-B Package-to-group Mechanism
The package-to-group mechanism bundles multiple packages with similar purposes into a group, as shown in Figure 1. This mechanism primarily simplifies the installation and management, enhancing efficiency for developers and maintainers. Different Linux distributions implement this mechanism in different ways. For instance, Debian and Debian-based Linux distributions (e.g., Ubuntu) use metapackages [30], which defines groups of multiple packages based on their purposes. This allows developers and maintainers to integrate a specific functional feature in one step, i.e., install a corresponding group with all relevant packages. In Red Hat and its derivatives (e.g., Fedora and CentOS), this mechanism is based on the concept of groups. These groups can be defined by a configuration file, which lists all groups and their associated packages. Maintainers can easily install or uninstall groups, making it easier to manage large sets of packages.
The P2G mechanism not only helps maintainers more easily locate and install the packages they need but also facilitates the maintenance of OSS projects. Additionally, this mechanism can simplify security updates and upgrades, as OSS projects can be targeted at specific groups rather than updating each package individually. This is particularly important in managing servers and large enterprise environments.
@@ III Overview
In this section, we present the design of our empirical study and provide an overview of the selected Linux distributions.
@@ III-A Study Design
Figure 2 provides an overview of our empirical study, which focuses on five Linux distributions (Section III-B). Our study follows the evolutionary process and can be divided into three phases (i.e., the past, the present, and the future). In the past phase, we conduct quantitative analysis and change analysis on historical versions to investigate the P2G’s application (RQ1) and evolution (RQ2). In the present phase, we assess the latest version’s quality (RQ3) by proposing metrics, evaluating groups, and summarizing issues. In the future phase, we conduct a topic analysis for group descriptions and key information extraction from package descriptions to explore group tendency (RQ4). The practical implications of our study include providing current Linux maintainers with update recommendations based on evolutionary patterns and quality issues and helping Linux developers determine whether to adopt P2G based on group tendency.
@@
III-B Studied Linux Distributions
There are many package management tools for Linux, such as RPM for Red Had-based distributions and DEB for Debian-based distributions. Since RPM has become one of the recognized standards in Linux distributions [12], we rank RPM-based distributions according to developers’ interests [31] and select the two most popular open-source distributions, i.e., Fedora and CentOS. They are widely used general-purpose distributions. To provide broader coverage of Linux across different scenarios, we include another three distributions tailored for specific cases, i.e., OpenEuler [zhou2022openeuler], Anolis [32] and OpenCloudOS [33]. OpenEuler is optimized for cloud and edge computing and is suitable for industry-specific applications. Anolis focus on enterprise-level applications and is a popular choice for data centers. OpenCloudOS is designed for could computing environments and is suitable for cloud service providers. Additionally, they are all representative distributions from different companies. Table I shows the studied Linux distributions and their versions in our study. We obtained package and group information for each distribution from official or widely available mirrors.
@@ IV-A Methodology
We answer this RQ from three aspects: 1) the trend in the number of groups, the number of packages adopting P2G, and the total number of packages; 2) the distribution of packages adopting P2G within a Linux distribution; 3) the relationship between the adoption of P2G and distribution popularity. The analysis is based on the selected Linux distributions and release versions in Section III-B. To decide whether a package adopts P2G in a specific release version, we consider whether it is included in a group in that version.
@@ IV-B Results
Application Trends. Figure 3 shows the number of groups, packages adopting P2G, and total packages across all historical release versions. ❶ We observe that the number of groups remains relatively stable whether in a complex Linux distribution (e.g., Fedora) or a relatively simplified distribution (e.g., OpenEuler). This suggests that Linux distributions have a relatively fixed group structure. ❷ However, most Linux distributions show an increase in the number of packages adopting P2G, especially in more innovative and developer-oriented distributions (e.g., Fedora). In contrast, in highly stable scenarios (e.g., enterprise or server environments), Linux distributions (e.g., openEuler) show less variation, as they prioritize reliability and compatibility over frequent changes. ❸ The number of total packages continues to increase with version updates. This reflects that Linux distributions continuously offer additional features and support in response to enhanced functionality and growing user requirements. For example, the number of packages in Fedora increased nearly tenfold, from 11,410 to 103,176.
@@
Distribution of packages adopting P2G. Figure 4 shows violin plots with the distribution and ratio of packages adopting P2G. ❶ The number of packages adopting P2G varies significantly across distributions. Fedora has a median of 3,544 packages adopting P2G and OpenEuler has only 776, which is approximately one-fifth of Fedora’s. ❷ The distribution of package numbers across versions also varies among Linux distributions. Anolis shows significant fluctuations, ranging from 0 to 4,000, while OpenEuler remains within a more stable range of 700 to 800, indicating minimal volatility. ❸ The number and ratio of packages adopting P2G are not consistent. Fedora has the highest number of grouped packages (median of 3,544), but a lower ratio of 0.27. In contrast, CentOS has fewer grouped packages (a median of 1,297), but it has the highest ratio at 0.47.
@@
Distribution of ratio per version
@@
Relationship between the popularity of Linux distributions and the ratio of packages adopting P2G. As shown in Figure 5, the ratio shows a downward trend, except for Anolis. It suggests that a Linux distribution’s popularity is correlated with the application of P2G. To validate this assumption, we perform a Spearman correlation analysis between the ratio of packages adopting P2G and the number of stars a Linux distribution has. However, the results (p-value=0.6, rho=0.4) indicate a weak positive correlation and the high p-value suggests that this correlation is statistically insignificant. Thus, it shows that the factors related to a Linux distribution’s popularity are complex. The anomaly observed in Anolis may be due to its inheritance of the CentOS ecosystem, resulting in a higher ratio despite its relatively low popularity.
Figure 5: Ratio of the packages applied P2G mechanism. All the Linux distributions are sorted in descending order of popularity (#stars).